### v1

# The HLLC-2D method for the computation of two phase flow system ejecta transporting model

Abstract
This work presents a second-order arbitrary Lagrangian Eulerian (ALE) method for the compressible multiphase particle-flow system with the HLLC-2D Riemann solver. We focus on researching the precise equation to describe the interactions between partible phase and flow phase. The computation of equivalent relation of the momentum and energy between two phases is the key point during the procedure, which can be capable of maintaining the conservation of this system. In the case of particle terms, track their trajectories within the mesh and couple with the element based on the position in the space. Thereafter an ALE method instead of Lagrangian scheme is derived for the discretization of the equation to perform better with the complex motion of particles and flow. We apply the HLLC-2D Riemann solver to substitute the HLLC solver which relaxes the requirement for continuous fluxes along the edge. Meanwhile we propose a method for searching particles and provide a CFL-like condition based on this. Finally, we give some numerical tests to analysis the influence of particles on fluid and get a following effect between two phases. The model and method are validated through numerical tests to show its robustness and accuracy.
Keywords:  Multiphase flow, ejecta model, ALE method, HLLC-2D
2010 Mathematics Subject Classification:   65M06, 76M12, 76T99
1 Introduction
The compressible multiphase flows in a gas-particle system is one of the most important problems in both nature and engineering systems. Due to the high-complexity and multi-scale natures of the gas-particle system, giving an experimental methods of gas-particle system is an investigation so that the numerical simulation of multiphase flows is necessary in the research of the gas-particle flow.
Generally in the early time, systematic research on two-phase flow began in the 1940s, but it was primarily focused on experimental observations and phenomenological descriptions. The term ’Two-phase Flow’ first appeared in the literature in [15]. After that, people started to focus on this problem which led to the gradual emergence of numerous research papers and monographs on multiphase flow theory. Over the past few years, the study of two-phase flow has evolved from relatively simple pipe flows to more complex fluidized beds and gas-particle two-phase separation flows. It has also progressed from low concentration (dilute particle phase) flows to high concentration (dense particle phase) flows. Building upon the foundation of dilute phase flow, researchers have developed physical models and fundamental equations for dense phase flow[22, 23, 24, 25, 26]. Subsequently, many researchers, including Campbell et al.[27], conducted a series of studies and established the theory of granular dynamics. Then Ding and Gidaspow[28] further developed and refined the theory of granular dynamics, introducing the concept of virtual temperature to characterize the effects of particle collisions in dense granular flows. They also performed numerical simulations that yielded results consistent with experimental observations.
In the recent years, the complex interactions in multiphase flow between phases have garnered increasing attention. Many scholars have conducted detailed research on turbulent disturbances in two-phase flows, with particular emphasis on simulating gas-solid two-phase turbulence, which has become a research focus. Various approaches have been developed, including tow-phase turbulence direct numerical simulation[29], Reynolds-averaged numerical simulation and large-eddy simulation[51, 50].
Especially in two-phase problems, subject to extreme shock loading, metal materials may undergo fragmentation, resulting in high-speed disintegration of particulate matter separating from the metal materials. This particulate matter then mixes with the surrounding gas, leading to a phenomenon called ’ejecat mixing’[34, 35]. The theoretical and experimental method and numerical simulation of ejecta mixing is currently a challenging and prominent issue in the field of implosion compression science [29, 30, 31, 33]. It represents one of the significant challenges in two-phase flow and is also a highly important area of application.
There are many experimental results publishing on the ejecta mixing. In 1983, Elias et al. published the experimental results of jet mixing involving metallic tin and argon gas [36]. They conducted joint observations of ejecta mixing evolution using X-ray imaging and high-speed photography. Later Ogorod.nikov et al. designed ejecta mixing experiments involving metallic lead and air[37]. In the same year, Buttler et al. [39] utilized PDV technology to investigate the velocity evolution of tin ejecta in neon gas. Their experimental results reflected the gradual decay of the ejecta over time due to fragmentation and resistance effects. In [38], Oro et al. conducted ejecta mixing experiments by preloading tungsten particles with a certain size distribution on an aluminum substrate using explosive loading, including vacuum ejecta mixing experiments, argon gas filling experiments, and hydrogen gas filling experiments. Sorenson et al. [40] further studied the characteristics of ejecta size distribution with different gases.
On the other hand, studying particle motion through numerical simulation is also an important role to supplement experiments.In 2004, Wang et al.[52] used SPH method to simulate the ejecta phenomenon on the metal surface under shock loading. Harrison, Fung, and others[11, 41], based on the FLAG Lagrangian fluid dynamics program platform, have developed an initial numerical simulation program for ejecta mixing. This program consists of two main modules: the particle source module and the particle transport module. The particle source module is used to generate and initialize particles, including determining the conditions for ejecta initiation, the total mass of ejecta, the mass and velocity distribution of the ejecta, the size distribution of the particles, the thermodynamic state of the particles, initial velocity, initial position, and more. The particle transport module is employed to simulate the movement of particles in the gas environment, with the main algorithm utilizing the MP-PIC method [44, 45, 12]. In 2013, Fung et al. [42] provided a detailed introduction to the method of introducing initial particles and their state settings, but did not present quantitative calculations of ejecta mixing results or comparisons with experimental data.
Due to the complexity of the coupled fluid-particle system, it means that while the fluid influences the particles, the particles also exert an effect on the fluid, necessitating the use of robust numerical methods for fluid model computations. In recent years, a classical numerical method for fluid equation is primarily based on a conserved cell-centered grid discretization. The one-dimensional scheme of this method was initially proposed by Godunov et al. [46, 13], and later it was extended to multi-dimensional scenarios [47, 48]. And later many scholars have proposed approximate Riemann solvers at the nodes, such as Despres and Mazera [49], Maire [3], which have shown significant improvements in correcting results.
Drawing from the advancements outlined in the above-mentioned literature, we introduce a coupled particle-fluid transporting model by configuring the initial states of the particles in a case of low concentration. And we focus on solving particle-gas system transporting problem based on a cell-centered Arbitrary Lagrangian-Eulerian (ALE) method which is the generalization and application from the ALE frame in [1].
Based on the ejecta mixing and a coupled transporting effect, we apply the particle trajectory model to analyze the interactions and energy exchange between gas phase and particle. Through the numerical tests, we show the motion results of particles and fluid and a following effect of particles in fluid. Furthermore, we analyze the parameters influencing this effect and the reasons for its formation.
The rest of the paper are organized as follows. In Section 2, we present the particle trajectory model and give the governing equations. In Section 3, a numerical simulation method is constructed. In Section 4 we give the time discretization and a particle searching strategy for an extra CFL-like condition to control the time step. In Section 5 we show the numerical tests, analyze the results and give a conclusion.
2 Particle Trajectory Model
In this section, we shall study the particle trajectory model by analyzing the motion of particle and fluid. It is a two-dimensional model describing the compressible two-phase problem. Though multiphase flow are more complex than single-phase flow, they still follow the basic of mass, momentum and energy conservation in essence. So the model must maintain the conservation of the system.
2.1 Problem statement, Notation and Assumptions
Next we will introduce an ejecta mixing model considering the following problem. Multiple particles mix in the fluid(the total volume is less than a certain percentage). Regard each particle as an individual, interacting with the fluid separately and explore the motion of each particle independently. Since our particle density is less than a certain range, the particle flow is considered to be sparse particle flow, so it is assumed that each particle does not affect each other. And the fluid has initial density, velocity, pressure and interacts with particles to exchange momentum and energy at the same time.
In the following part, subscript $p$ indicates the physical parameter that is about particle phase, subscript $g$ is about the gas phase. Let $\rho_{g}$, $\bm{v_{g}}=(v_{g_{x}},v_{g_{y}})$, $E_{g}$, $\mu_{g}$, $T_{g}$, $C_{g}$ be the density, velocity, energy, viscosity and thermal conductivity of the fluid and $P$ be the pressure of the space. $\bm{u_{p}}=(x,y)$, $\rho_{p}$, $\bm{v_{p}}=(v_{p_{x}},v_{p_{y}})$, $T_{p}$, $c_{p}$, $r_{p}$ denote the coordinate, density, velocity ,temperature, specific heat capacity and radius of particle. In particular, $\bm{u_{p_{m}}}=(x_{m},y_{m})$, $\rho_{p_{m}}$, $\bm{v_{p_{m}}}$, $T_{p_{m}}$, $c_{p_{m}}$ $r_{p_{m}}$denote the variable of the particle $m$, in which $m=1,\ldots,m(p)$, $m(p)$ is the total number of particles.
To explore the interaction between two phases, the force analysis of the particles in the system is required. During the movement, particles are subject to gravity, drag, differential pressure, Basset force, Magnus force, Saffman force, etc. In general, not all of these forces are equally important. It is necessary to compare the magnitude of the above forces in order to estimate the relative importance of the various forces. Under normal conditions, other forces are several orders of magnitude smaller than the drag. So for the convenience of analysis, above forces except drag are not considered in the following model.
In addition to the force analysis, the transfer of heat between the particles and the fluid is also significant. Normally there are three ways for transferring the heat including heat conduction, convection heat transfer, thermal radiation. Similarly, because of the orders of the magnitude, we only consider the convection heat transfer.
To sum up, give the following assumptions:
  • (1)
    Consider ideal compressible gas. For each case we use an ideal gas equation of state, namely $P=(\gamma-1)\rho\epsilon$ , where $\gamma$ is the adabatic index.
  • (2)
    Viscosity and heat transfer exist only in interphase interactions.
  • (3)
    The motion of particles only take into account the drag due to friction.
  • (4)
    The particles are ideal round or spherical, uniform in diameter.
  • (5)
    There is no collision and interaction between particles.
  • (6)
    All the energy produced by friction is absorbed by the fluid.
2.2 Equations
According to the above assumptions, we introduce the equation of the system.
2.2.1 Governing equation
The governing equations for two-dimensional compressible particle trajectory model are as follows:
$\frac{\partial\rho_{g}}{\partial t}+\nabla\cdot(\rho_{g}\bm{v}_{g})=0,$ (2.1)
$\frac{\partial}{\partial t}(\rho_{g}\bm{v}_{g})+\nabla\cdot(\rho_{g}\bm{v}_{g}\otimes\bm{v}_{g})+\nabla P=-\bm{F}_{p},$
$\frac{\partial}{\partial t}(\rho_{g}E_{g})+\nabla\cdot(\rho_{g}E_{g}\bm{v}_{g})+\nabla\cdot(p\bm{v}_{g})=-{E_{p}},$
$\frac{\partial\bm{u_{p}}}{\partial t}=\bm{v}_{p},$ (2.2)
$\frac{\partial\bm{v}_{p}}{\partial t}=\bm{F}_{p},$
$\frac{4}{3}\pi\,r_{p}^{3}\,\rho_{p}\,c_{p}\frac{\partial T_{p}}{\partial t}=Q_{p},$
where the $\bm{F}_{p}$ and ${E_{p}}$ indicate the exchange of momentum and energy due to interaction between the particles and fluid. In the particle trajectory model, the motion of every particles will be calculated respectively, we have
$\bm{F}_{p}=\sum_{m=1}^{m(p)}\bm{F}_{p_{m}},\$ (2.3)
${E_{p}}=\sum_{m=1}^{m(p)}{E}_{p_{m}}.\$ (2.4)
$\bm{F}_{p_{m}}$ is the force between particle ${}^{\prime}m^{\prime}$ and the fluid which is the source power of the particle’s movement. ${E}_{p_{m}}$ is the exchange of energy including the heat and the kinetic energy between the particle ${}^{\prime}m^{\prime}$ and fluid.
2.2.2 Interaction and Conservation
Next we give the specific definition of interaction items $\bm{F}_{p}$ and ${E_{p}}$.
The force $\bm{F}_{p}$ comes from the drag due to friction. There are two drag models usually used during computing which are the Stoke’s model and Crowe’s model.
For the Stoke’s model[5], the following formulation is applied:
$\bm{F}_{p}=\left({F}_{p_{1}},{F}_{p_{2}}\right)^{T}=6\pi r_{p}\mu_{g}(\bm{v}_{g}-\bm{v}_{p}),$ (2.5)
where $r_{p}$ is the radius of the particles and $\mu_{g}$ is the gas viscosity.
For the Crowe’s model[6], we have
$\bm{F}_{p}=\frac{1}{2}C_{d}S_{p}\rho_{g}\left|\bm{v}_{g}-\bm{v}_{p}\right|\left(\bm{v}_{g}-\bm{v}_{p}\right),$ (2.6)
where $C_{D}$ is the drag coefficient and $S_{p}$ is the upwind area. Drag coefficient $C_{D}$ is related to the Reynolds number $Re$, the following definition is applied[4][8]:
$C_{D}=\frac{24}{Re_{p}}f_{drag}(Re).$ (2.7)
Function $f_{drag}(Re)$ depends on the Reynolds number $Re$, it has:
$Re=\frac{2\rho_{g}r_{p}\left|\bm{v}_{g}-\bm{v}_{p}\right|}{\mu_{g}},$
and
$C_{\mathrm{D}}=\begin{cases}\frac{24}{Re},&Re<0.2,\\ \frac{24}{Re}\left(1+0.15Re^{0.687}\right),&0.2\leqslant Re\leqslant 800,\\ 0.5,&Re>800.\\ \end{cases}$ (2.8)
Remark: The calculation method of $C_{\mathrm{D}}$ is not same in different research. The above definition is given by Wang et al in paper [8].
The energy interaction consists of two parts of which are drag work and heat transfer, namely:
${E_{p}}=\bm{F}_{p}\cdot\bm{v}_{p}+Q_{p}.$ (2.9)
$Q_{p}$ comes from the convection heat transfer, which is[7]:
$Q_{p}=\frac{2\mu_{g}}{Pr}c_{pg}\pi r_{p}\left(T_{g}-T_{p}\right)Nu_{p},$ (2.10)
where $T_{g}$ and $T_{p}$ is the temperature of fluid and particle respectively. $C_{pg}$ is constant-pressure specific heat capacity, $Nu_{p}$ is Nusselt numbers, $Pr$ is Prandtl numbers of flow, of which specific definitions are as follows:
$Pr=\frac{4\gamma}{9\gamma-5},$ (2.11)
and
$Nu_{p}=2+0.6Pr^{1/3}Re^{1/2}.$ (2.12)
Based on the equation (2.1) and (2.2), we need to prove that the above gas-particle two phase system satisfies conservation law. It is obvious that the conservation of momentum because the drag between the particle and the fluid is a pair of interacting forces, namely the right items of the second equation of (2.1) and of the equation (2.2). The item $\bm{F}_{p}\bm{v}_{p}$ is the energy of the fluid which is converted into the kinetic energy of the particle. The convection heat transfer energy $Q_{p}$ is generated by the temperature difference between the fluid and the particle, which is from the higher phase to the lower phase.
3 Numerical Scheme
In this section,we will give a numerical scheme to solve the equation (2.1) and (2.2).
3.1 ALE scheme
For the convenience, recast the equation (2.1) as follows:
$\frac{\partial\mathbf{U}}{\partial t}+\frac{\partial\mathbf{F}(\mathbf{U})}{\partial x}+\frac{\partial\mathbf{G}(\mathbf{U})}{\partial y}=\mathbf{S}_{p},$ (3.1)
where the state vector and flux vectors are
$\mathbf{U}=\left[\begin{array}[]{c}\rho_{g}\\ \rho_{g}v_{g_{x}}\\ \rho_{g}v_{g_{y}}\\ \rho_{g}E_{g}\end{array}\right],\quad\mathbf{F}(\mathbf{U})=\left[\begin{array}[]{c}\rho_{g}v_{g_{x}}\\ P+\rho_{g}v_{g_{x}}^{2}\\ \rho_{g}v_{g_{x}}v_{g_{y}}\\ \rho_{g}E_{g}v_{g_{x}}+Pv_{g_{x}}\end{array}\right],\quad\mathbf{G}(\mathbf{U})=\left[\begin{array}[]{c}\rho_{g}v_{g_{y}}\\ \rho_{g}v_{g_{x}}v_{g_{y}}\\ \rho_{g}v_{g_{y}}^{2}+P\\ \rho_{g}E_{g}v_{g_{y}}+Pv_{g_{y}}\end{array}\right],$ (3.2)
and $\mathbf{S}_{p}$ is the resource term between the flow and particles
$\mathbf{S}_{p}=\left[\begin{array}[]{c}0\\ {F}_{p_{1}}\\ {F}_{p_{2}}\\ {E}_{p}\end{array}\right].$ (3.3)
Based on ALE method, we give the numerical scheme. For the subsequent discretization process, it is advantageous to express equation (2.1) in the form of a moving control volume formulation:
[…]
$\mathbf{w}$ represents the moving velocity of the control volume $\Omega$ while $\mathbf{N}$ denotes the unit outward normal direction along the boundary of $\Omega$. When $\mathbf{w}=\mathbf{u}$, the system simplifies into a Lagrangian formulation, and when $\mathbf{w}=\mathbf{0}$, it takes on an Eulerian form.
Remark 3.1.
It is important to note that in the equation (3.4) the integration domain $\Omega$ is related to time $t$, that is $\Omega=\Omega\left(t\right)$. For convenience, we denote it as $\Omega$.
3.1.1 Notations on a generic polygonal grid
Every cell $\Omega_{c}$ within the mesh is assigned a distinct index represented by $c$. To refer to an adjacent cell $\Omega_{d}$ that shares a common edge with $\Omega_{c}$, we employ the subscript $d$. The edge where two units intersect is defined as $c\cap d$ or denoted as $k$, as shown in Fig 2. ${F}(c)$ represents the collection of neighboring cells surrounding cell $\Omega_{c}$. and ${Q}(c)$ is the set of all vertices of a cell $\Omega_{c}$. Each vertex within the mesh is allocated a unique index $q$, and $C(q)$ and $K(q)$ are defined as the sets of cells and edges respectively surrounding the vertex $q$, i.e.,
$C(q)=\left\{\Omega_{c}\text{ : cells sharing a vertex }q\right\},$
$K(q)=\{k\text{ : edges sharing a vertex }q\}.$
The physical quantities of flow such as the density $\rho_{g_{c}}$, pressure $P_{g_{c}}$, velocity $\bm{v}_{g_{c}}$, energy $E_{g_{c}},e_{g_{c}}$ are defined in the cell center of $\Omega_{c}$. The grid moving velocity, denoted as $\mathbf{w}_{q}$, is defined on the node $q$. Let $\mathbf{N}_{d}^{c}=\left(\left(n_{x}\right)_{d}^{c},\left(n_{y}\right)_{d}^{c}\right)$ represents the unit normal vector of cell $\Omega_{c}$ along edge $c\cap d$, with the subscript $d$ and the superscript $c$ indicating that the vector direction originates from $\Omega_{c}$ and extends towards $\Omega_{d}$.
Figure 1: Notations on a grid
Based on the edge’s normal direction between two adjacent cells, we utilize indices $L$ and $R$ to represent the states on the left and right sides of the edge, respectively. For example, $\Omega_{c}$ and $\Omega_{d}$ are the left and right cells of the edge $c\cap d$.
To compute integration of the variable $\mathbf{F}$ across face $\left[M_{a},M_{a+1}\right]$, denote by $\mathbf{F}_{a,a+1}$ the fluxes on the face, $L_{a,a+1}$ is the length of $\left[M_{a},M_{a+1}\right]$, we have
$L_{a,a+1}\mathbf{F}_{a,a+1}\cdot\mathbf{N}_{\bm{a},\bm{a}+\mathbf{1}}=\int_{M_{a}}^{M_{a+1}}\mathbf{F}\cdot\mathbf{N}\mathrm{d}l.$ (3.5)
Using the above transformation to discretize the equations (3.4) by utilizing the idea of the Godunov scheme on unstructured mesh. Define cell-centered state variable $\mathbf{U}_{c}$ and resource term $\mathbf{S}_{p_{c}}$, then we have:
[…]
$={\left|\Omega_{c}\right|}\mathbf{S}_{p_{c}}-\sum_{d}L_{d}^{c}\left(\mathcal{T}_{d}^{c}\right)^{-1}\mathbf{F}_{d}^{c}\left(\mathbf{w}_{d}^{c}\cdot\mathbf{N}_{d}^{c},\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right),$
where $\left|\Omega_{c}\right|$ represent the volume of cell $\Omega_{c}$. $\mathbf{w}_{d}^{c}$ is the moving velocity which is defined at the center of cell edge $c\cap d$. $L_{d}^{c}$ is the length of the edge $c\cap d$. $\mathcal{T}_{d}^{c}$ and $\left(\mathcal{T}_{d}^{c}\right)^{-1}$ are the rotation matrix and its inverse as follows:
$\mathcal{T}_{d}^{c}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\left(n_{x}\right)_{d}^{c}&\left(n_{y}\right)_{d}^{c}&0\\ 0&-\left(n_{y}\right)_{d}^{c}&\left(n_{x}\right)_{d}^{c}&0\\ 0&0&0&1\end{array}\right),\quad\left(\mathcal{T}_{d}^{c}\right)^{-1}=\left(\begin{array}[]{cccc}1&0&0&0\\ 0&\left(n_{x}\right)_{d}^{c}&-\left(n_{y}\right)_{d}^{c}&0\\ 0&\left(n_{y}\right)_{d}^{c}&\left(n_{x}\right)_{d}^{c}&0\\ 0&0&0&1\end{array}\right).$ (3.7)
There are many methods to calculate the approximate numerical flux $\mathbf{F}_{d}^{c}$ between the cell $\Omega_{c}$ and $\Omega_{d}$. The most commonly used involves solving a one-dimensional Riemann problem along the normal direction of edge $c\cap d$ :
$\mathbf{U}_{t}+\mathbf{F}_{x}=\mathbf{S}_{p}$ (3.8)
$\mathbf{U}(x,0)=\begin{cases}\mathbf{U}_{L}=\mathcal{T}_{d}^{c}\mathbf{U}_{c},&x<0,\\ \mathbf{U}_{R}=\mathcal{T}_{d}^{c}\mathbf{U}_{d},&x>0,\end{cases}$
and the numerical flux can be represented as follows:
$\mathbf{F}_{d}^{c}\left(w,\mathbf{U}_{L},\mathbf{U}_{R}\right)=\mathbf{F}^{w}-w\mathbf{U}^{w},$ (3.9)
where $w=\mathbf{w}_{d}^{c}\cdot\mathbf{N}_{d}^{c}$ is the normal projection of the moving velocity $\mathbf{w}_{d}^{c}$ , $\mathbf{U}^{w}$ and $\mathbf{F}^{w}$ are the projection of the state and flux functions onto the direction $w=x/t$ respectively.
3.1.2 Flux formula of HLLC-2D on moving mesh
We use the two-dimensional Riemann solver HLLC-2D method proposed in [2, 1] to evaluate the numerical flux. The construction of this methods hinges on two important points. Firstly, it entails a nodal solver that relies on the vertex $q$ with multi-state values $\left\{\mathbf{U}_{c},c\in C(q)\right\}$ assigned to neighbor cells. Then the nodal contact velocity $\mathbf{u}_{q}^{*}$ is determined through this solver. Secondly, it involves the utilization of the HLLC-2D solver, which is an expanded scheme of the classical one-dimensional HLLC Riemann solver. In Riemann problem (3.8), we have the nodal flow velocity $\mathbf{u}_{q}^{*}$, as well as the state vectors $\mathbf{U}_{c}$ and $\mathbf{U}_{d}$, which are on the sides of edge $c\cap d$. Through solving the Riemann problem, we could get the approximate numerical flux $\mathbf{F}_{c}^{d}$ and $\mathbf{F}_{d}^{c}$ and further solving the equation (3.6).
In the following evaluation, assume that the nodal contact velocity $\mathbf{u}_{q}^{*}$ has already been determined, in alignment with the preceding chapter’s description. From the equation (3.8), the initial values of the equation are:
$\mathbf{U}_{L}=\mathcal{T}_{d}^{c}\mathbf{U}_{c},\quad\mathbf{U}_{R}=\mathcal{T}_{d}^{c}\mathbf{U}_{d},$
so that the corresponding fluxes are $\mathbf{F}_{L}=\mathbf{F}\left(\mathbf{U}_{L}\right),\mathbf{F}_{R}=\mathbf{F}\left(\mathbf{U}_{R}\right)$.
In contrast to the classical one-dimensional Riemann solver, HLLC-2D method deviates by introducing an artificial contact wave speed denoted as $S_{*}=\mathbf{u}_{q}^{*}\cdot\mathbf{N}_{d}^{c}$, which is consistent with the nodal velocity $\mathbf{u}_{q}^{*}$.
Establish two velocities $S_{L}$ and $S_{R}$ and suppose $S_{L}\leqslant S_{*}\leqslant S_{R}$. As a result, the HLLC-2D solver comprises 4 constant state regions separated by $S_{L},S_{*}$ and $S_{R}$. Within these regions, the vector of conservation quantities is defined as follows:
$\mathbf{U}^{w}=\begin{cases}\mathbf{U}_{L},&\text{ if }w\leqslant S_{L},\\ \mathbf{U}_{L}^{*},&\text{ if }S_{L}<w\leqslant S_{*},\\ \mathbf{U}_{R}^{*},&\text{ if }S_{*}<w\leqslant S_{R},\\ \mathbf{U}_{R},&\text{ if }w>S_{R},\end{cases}$ (3.10)
where $\mathbf{U}_{L}^{*},\mathbf{U}_{R}^{*}$ are the same as the classical HLLC solver. Within each state region, we introduce the two approximate fluxes which is from the different direction of the edge $c\cap d$:
$\mathbf{F}_{H}^{w}=\begin{cases}\mathbf{F}_{H,1},&\text{ if }w\leqslant S_{L},\\ \mathbf{F}_{H,2},&\text{ if }S_{L}<w\leqslant S_{*},\\ \mathbf{F}_{H,3},&\text{ if }S_{*}<w\leqslant S_{R},\\ \mathbf{F}_{H,4},&\text{ if }w>S_{R},\end{cases}$ (3.11)
where $H=L,R$, and the above flux scheme ensure strictly satisfies the R-H conditions.
According to the equation (3.9) and flux struction (3.11). Numerical fluxes encircling the vertex q on the moving mesh are as follows:
$\mathbf{F}_{d,q}^{c}\left(w,\mathbf{U}_{L},\mathbf{U}_{R}\right)=\mathbf{F}_{L}^{w}-w\mathbf{U}^{w}=\begin{cases}\mathbf{U}_{L}\left(u_{L}-w\right)+\mathbf{D}_{L},&\text{ if }w\leqslant S_{L},\\ \mathbf{U}_{L}^{*}\left(S_{*}-w\right)+\mathbf{D}_{L}^{*},&\text{ if }S_{L}<w\leqslant S_{*},\\ \mathbf{U}_{R}^{*}\left(S_{*}-w\right)+\mathbf{D}_{L}^{*},&\text{ if }S_{*}<w\leqslant S_{R},\\ \mathbf{U}_{R}\left(u_{R}-w\right)+\mathbf{D}_{R}-\mathbf{D}_{R}^{*}+\mathbf{D}_{L}^{*},&\text{ if }w>S_{R},\end{cases}$ (3.12)
and
$\mathbf{F}_{c,q}^{d}\left(w,\mathbf{U}_{L},\mathbf{U}_{R}\right)=\mathbf{F}_{R}^{w}-w\mathbf{U}^{w}=\begin{cases}\mathbf{U}_{L}\left(u_{L}-w\right)+\mathbf{D}_{L}-\mathbf{D}_{L}^{*}+\mathbf{D}_{R}^{*},&\text{ if }w<S_{L},\\ \mathbf{U}_{L}^{*}\left(S_{*}-w\right)+\mathbf{D}_{R}^{*},&\text{ if }S_{L}\leqslant w<S_{*},\\ \mathbf{U}_{R}^{*}\left(S_{*}-w\right)+\mathbf{D}_{R}^{*},&\text{ if }S_{*}\leqslant w<S_{R},\\ \mathbf{U}_{R}\left(u_{R}-w\right)+\mathbf{D}_{R},&\text{ if }w\geqslant S_{R},\end{cases}$ (3.13)
where $\mathbf{D}_{H}=(0,P,0,Pu)^{T},\mathbf{D}_{H}^{*}=\left(0,P_{H}^{*},0,P_{H}^{*}S_{*}\right)^{T}$ for $H=L,R$ are vectors including the pressure and work terms on ends and in the middle two regions, respectively. Pressures are
$\left\{\begin{array}[]{l}P_{L}^{*}=p_{L}+\rho_{L}\left(S_{L}-u_{L}\right)\left(S_{*}-u_{L}\right),\\ P_{R}^{*}=p_{R}+\rho_{R}\left(S_{R}-u_{R}\right)\left(S_{*}-u_{R}\right).\end{array}\right.$ (3.14)
An analogous derivation can be given at vertex $q^{+}$, and flux $\mathbf{F}_{d,q^{+}}^{c}$ is evaluated. For the edge $c\cap d$, the flux on it finally is:
$\mathbf{F}_{d}^{c}\left(w,\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)=\frac{1}{2}\left(\mathbf{F}_{d,q}^{c}\left(\mathbf{w}_{q}\cdot\mathbf{N}_{d}^{c},\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)+\mathbf{F}_{d,q^{+}}^{c}\left(\mathbf{w}_{q^{+}}\cdot\mathbf{N}_{d}^{c},\mathcal{T}_{d}^{c}\mathbf{U}_{c},\mathcal{T}_{d}^{c}\mathbf{U}_{d}\right)\right)$ (3.15)
3.1.3 Nodal Solver
In Subsection 3.1.2, we have assumed that the nodal contact velocity is already calculated and then present a nodal solver that consistently evaluate the $\mathbf{u}_{q}^{*}$ to ensure that the scheme maintains local conservation of mass, momentum and total energy. Since $\mathbf{F}_{d}^{c}\neq\mathbf{F}_{c}^{d}$ may happen, the conservation in cells is not satisfied. Refer to the construction of the solver, we want to ensure the conservation of mass, momentum and total energy enforced at nodes and we could derive this procedure from the global conservation within domain $\Omega$. By maintaining the local nodal conservation including mass, momentum and total energy, contact velocity could be calculated. By using the pressure equation (3.14), we have:
$\sum_{k\in K(q)}L_{k}\left(\alpha_{L,k}+\alpha_{R,k}\right)\left[\mathbf{u}_{q}^{*}\cdot\mathbf{N}_{k}-v_{k}^{*}\right]\mathbf{N}_{k}=0,$ (3.16)
where
$\alpha_{L,k}=-\rho_{L,k}\left(S_{L,k}-u_{L,k}\right),\quad\alpha_{R,k}=\rho_{R,k}\left(S_{R,k}-u_{R,k}\right),$ (3.17)
and $v_{k}^{*}$ is the contact velocity in the classical one-dimensional HLLC Riemann solver for edge $k$
$v_{k}^{*}=\frac{P_{L,k}-P_{R,k}+\alpha_{L,k}u_{L,k}+\alpha_{R,k}u_{R,k}}{\alpha_{L,k}+\alpha_{R,k}}.$ (3.18)
Solve the equation (3.16) to calculate the contact velocity $\mathbf{u}_{q}^{*}$ at each vertex $q$. Then we get
$\mathbf{u}_{q}^{*}=M^{-1}\sum_{k\in K(q)}L_{k}\left(\alpha_{L,k}+\alpha_{R,k}\right)v_{k}^{*}\mathbf{N}_{k},$ (3.19)
where
$M=\left[\begin{array}[]{cc}\sum_{k\in K(q)}L_{k}\left(\alpha_{L,k}+\alpha_{R,k}\right)n_{x,k}^{2}&\sum_{k\in K(q)}L_{k}\left(\alpha_{L,k}+\alpha_{R,k}\right)n_{x,k}n_{y,k}\\ \sum_{k\in K(q)}L_{k}\left(\alpha_{L,k}+\alpha_{R,k}\right)n_{x,k}n_{y,k}&\sum_{k\in K(q)}L_{k}\left(\alpha_{L,k}+\alpha_{R,k}\right)n_{y,k}^{2}\end{array}\right].$
3.1.4 Boundary Condition
In this ALE method, it employs the convensional zero-order extrapolation method to enforce boundary conditions. For velocity-based boundaries, such as solid wall or piston, two distinct cases arise (1) When three boundary nodes are not collinear, we apply a straightforward extrapolation method. (2) In cases where three boundary nodes are collinear, we employ t mirror extrapolation which entails the construction of a symmetrical point for each vertex connecting a boundary node. This method ensures the same scheme (3.19) as used for internal vertex when calculating $\mathbf{u}_{q}^{*}$.
4 Time discretization
In this section, we discretize in time for the system (3.6) that describes the evolution of the variables in the cell $\Omega_{i}$. Suppose that there are $m(i)$ particles in cell $\Omega_{i}$, of which the index is $m=1,\ldots,m(i)$, then we use a second-order Runge-Kutta scheme:
$\hat{\mathbf{U}}_{i}^{n+1}$ $=\frac{\Delta t^{n}}{\left|\Omega_{i}^{n+1}\right|}\sum_{m}\mathbf{S}_{p_{i}}^{n}+\frac{\left|\Omega_{i}^{n}\right|}{\left|\Omega_{i}^{n+1}\right|}\mathbf{U}_{i}^{n}-\frac{\Delta t^{n}}{\left|\Omega_{i}^{n+1}\right|}[\sum_{d}L_{d}^{i}\left(\mathcal{T}_{d}^{i}\right)^{-1}\mathbf{F}_{d}^{i}\left(\mathbf{w}_{d}^{i}\cdot\mathbf{N}_{d}^{i},\mathcal{T}_{d}^{i}\mathbf{U}_{i}^{n},\mathcal{T}_{d}^{i}\mathbf{U}_{d}^{n}\right)],$ (4.1)
$\overline{\mathbf{U}}_{i}^{n}=\frac{1}{2}({\mathbf{U}}_{i}^{n}+\hat{\mathbf{U}}_{i}^{n+1})$ (4.2)
$\mathbf{U}_{i}^{n+1}=\frac{\Delta t^{n}}{\left|\Omega_{i}^{n+1}\right|}\sum_{m}\mathbf{S}_{p_{i}}^{n}+\frac{\left|\Omega_{i}^{n}\right|}{\left|\Omega_{i}^{n+1}\right|}\mathbf{U}_{i}^{n}-\frac{\Delta t^{n}}{\left|\Omega_{i}^{n+1}\right|}\sum_{d}L_{d}^{i}\left(\mathcal{T}_{d}^{i}\right)^{-1}\mathbf{F}_{d}^{i}\left(\mathbf{w}_{d}^{i}\cdot\mathbf{N}_{d}^{i},\mathcal{T}_{d}^{i}\overline{\mathbf{U}}_{i}^{n},\mathcal{T}_{d}^{i}\overline{\mathbf{U}}_{d}^{n}\right),$ (4.3)
and similarly
$\hat{\bm{v}}_{p_{m}}^{n+1}=\bm{v}_{p_{m}}^{n}+\frac{\Delta t^{n}}{\pi r_{p_{m}}^{2}{\rho}_{p_{m}}}\,\bm{F}_{p_{m}}^{n},$ (4.4)
$\hat{\bm{F}}_{p_{m}}^{n+1}={\bm{F}}_{p_{m}}(\frac{1}{2}(\hat{\bm{v}}_{p_{m}}^{n+1}+{\bm{v}}_{p_{m}}^{n})),$
$\hat{T}_{p_{m}}^{n+1}=T_{p_{m}}^{n}+{\Delta t^{n}}\,\frac{3}{4\pi r_{p_{m}}^{3}\rho_{p_{m}}c_{p_{m}}}Q_{p_{m}}^{n},$
$\hat{Q}_{p_{m}}^{n+1}={Q}_{p_{m}}(\frac{1}{2}({T}_{p_{m}}^{n}+\hat{T}_{p_{m}}^{n+1}))$
so that
$\bm{v}_{p_{m}}^{n+1}=\bm{v}_{p_{m}}^{n}+\frac{\Delta t^{n}}{\pi r_{p_{m}}^{2}{\rho}_{p_{m}}}\,\hat{\bm{F}}_{p_{m}}^{n+1},$ (4.5)
$\bm{u}_{p_{m}}^{n+1}=\bm{u}_{p_{m}}^{n}+{\Delta t^{n}}\,\hat{\bm{v}}_{p_{m}}^{n+1},$
$T_{p_{m}}^{n+1}=T_{p_{m}}^{n}+\Delta t^{n}\,\frac{3}{4\pi r_{p_{m}}^{3}\rho_{p_{m}}c_{p_{m}}}\hat{Q}_{p_{m}}^{n+1},$
4.1 Particle searching
Before computing the time step $\Delta t^{n}=t_{n+1}-t_{n}$, we need to give the process of searching the particles during a time step. In the numerical simulation, the interaction between the particles and flow is calculated in every cells instead of the whole space. That is, if consider the cell $\Omega_{i}$, the momentum and energy exchanged between the two phase is that all the particles in cell $\Omega_{i}$ interact with the fluid in $\Omega_{i}$. So before compute the physical properties we should know which particles are in the cell. At the beginning of time, the coordinates of particles are given. And after every time steps, we should search the particles and confirm the location of every particles. In the numerical simulations, we use the following method for searching the particles:
  • (1)
    Suppose that we know the coordinate of particle m at time $t_{n}$, $\bm{u_{m}^{n}}=(x_{m}^{n},y_{m}^{n})$. And the particle m is in the cell $\Omega_{i}$.
  • (2)
    Calculate the equation (4.5) and get the coordinate of particle m, $\bm{u_{m}^{n+1}}=(x_{m}^{n+1},y_{m}^{n+1})$ , at time $t_{n+1}$.
  • (3)
    Search the particle m in the updated geometrical characteristics of the cell. First determine whether the particle m is in the cell $\Omega_{i}$. If then, proceed to step 5.
  • (4)
    If not, locate the surrounding cells of $\Omega_{i}$, suppose the index of these cells are $c=1,\dots,n(i)$. Search the particle m in the cell $\Omega_{c}$, $c=1,\dots,n(i)$.
  • (5)
    Update the location of particle m.
Using the above method, we could get the updated location of particle k under the condition that we give a limitation to the time step $\Delta t^{n}$.
Figure 2: Particle searching
4.2 Time step limitation
Because of using the explicit Euler method, we need to observe some limitations on time step. They are as follows:
  • (1)
    The variation of cell volume is coherent with the vertex motion.
  • (2)
    A CFL-like condition to be a priori satisfied for entropy inequality.
  • (3)
    Make sure that the particles can be searched.
Suppose that we know the physical properties in the cell $\Omega_{i}$ and the geometrical characteristics of the cell $\Omega_{i}$ at the beginning of the time step $t_{n}$. We will compute the variables’ value at time $t_{n+1}$ and time step $\Delta t^{n}$.
4.2.1 CFL condition
First we derive a CFL condition to ensure a positive entropy production in cell $\Omega_{i}$ during the time step. At time $t_{n}$, for each cell $\Omega_{i}$ we denote by $\lambda_{i}^{n}$ the inner circle radius of the cell. We have
$\Delta t_{e}=C_{e}\min_{i=1,\ldots,I}\frac{\lambda_{i}^{n}}{\sqrt{(u_{g_{i}}^{n})^{2}+(v_{g_{i}}^{n})^{2}}+c_{i}^{n}},$ (4.6)
where $C_{E}$ is a strictly positive coefficient and $c_{i}$ is the sound speed in the cell. The coefficient $C_{E}$ is given artificially. According to the numerical experiments ,$C_{E}=0.3$ is a value which could get good numerical results.
4.2.2 Criterion on the variation of volume
We also give the criterion on the variation of the cell volume. At time $t_{n}$, we can get the contact velocity of vertex, we have:
$\Delta t_{v}=C_{V}\min_{i=1,\ldots,I}\left\{\frac{\left|A_{i}^{n}\right|}{\left|A_{i}^{\prime}\right|}\right\}.\$ (4.7)
where $C_{v}=0.1$ and
$\left|A_{i}^{\prime}\right|=\frac{1}{2}\sum_{f}L_{f}^{c}\left(\mathbf{u}_{q}^{*}+\mathbf{u}_{q^{+}}^{*}\right)\cdot\mathbf{N}_{f}^{c}.$ (4.8)
4.2.3 Particle searching limitation
Refer to Subsection 4.1, suppose that particle $m$ is located in $\Omega_{m_{i}}$, we check the cell $\Omega_{m_{i}}$ and its surrounding cells. From time $t_{n}$ to $t_{n+1}$ the particles should be specify the range of motion of the particle. In the subsection 4.2.1 denote by $\lambda_{m_{i}}^{n}$ the inner circle radius of the cell. We have:
$\Delta t_{p}=C_{p}\min_{m}\min_{i=1,\ldots,n(m)}\frac{\lambda_{m_{i}}^{n}}{|\bm{v_{p_{m}}^{n}}|}.\$ (4.9)
where $n(m)$ is the number of cells that is containing and surrounding the particle. In numerical simulation we take $C_{p}=0.5$.
Finally, the estimation of the next time step the estimation of the next time step $\Delta t^{n}=t_{n+1}-t_{n}$ is given by
$\Delta t^{n}=\min\left(\Delta t_{e},\Delta t_{v},\Delta t_{p},1.01*\Delta t^{n-1}\right),\$ (4.10)
4.3 Algorithm for discretization
Give the description of the whole computation process:
Step 1. Initialization
At time $t=t^{n}$ we know in each cell $\Omega_{I},i=1,\ldots,I$, the fluid variables: $\bm{v}_{p_{i}}^{n}$, $E_{p_{i}}^{n}$, $\rho_{p_{i}}^{n}$, $c_{p_{i}}^{n}$, $P_{i}^{n}$, $\mu_{p_{i}}^{n}$,the particle variables: $\bm{v_{g_{i}}^{n}},x_{k_{i}}^{n},y_{k_{i}}^{n}$ and the geometrical quantities: $L_{k}^{n}$, $\mathbf{N}_{k}^{n}$ for $k\in K(q)$.
Step 2. Nodal solver of flow
The contact velocity $\bm{u}_{q}^{*}$ at each vertex is obtained by solving the linear system of equations in subsection 3.1.3.
Step 3. Time step limitation
Obtain the time stem limitation $\Delta t^{n}$ according to the (4.10).
Step 4. Grid moving strategy
In the ALE method, specific grid moving strategy, such as harmonic mapping equation or interpolation technique, needs to be used to determine vertex coordinate. Hence the nodal motion velocity $\mathbf{w}_{q}=\left(\mathbf{x}_{q}^{n+1}-\mathbf{x}_{q}^{n}\right)/\Delta t^{n}$. Futhermore, the cell volume could be obtained from the vertex coodinates.
Step 5. Interaction between two phase
Calculate the interaction $\mathbf{F}_{p}$ and ${E_{p}}$ at time $t_{n}$.
Step 6. The Riemann solver HLLC-2d
At the common edge $i\cap d$ between the cell $\Omega_{i}$ and $\Omega_{d}$, the edge fluxes $\mathbf{F}_{d,q}^{i}$ and $\mathbf{F}_{d,q^{+}}^{i}$ are calculated. The edge flux $\mathbf{F}_{d}^{i}$ is the average of the two half edge fluxes.
Step 7. Update the state variable of flow
Update the conservation quantities in cell $\Omega_{i}$.
Step 8. Update the particle
Calculate the variables of particles by equation (4.5) and search the location of particle at time $t_{n+1}$. Then it goes to Step 1 if the calculation continues, otherwise it stops.
5 Nemurical tests and results
In this section, we present several examples in order to validate our numerical scheme. We begin by several classical one-dimensional particle transporting tests in quiescent fluid, in a fluid under constant acceleration and in a fluid under sinusoidal acceleration. The relationship between particle size and particle motion path is further explored in this part. Next we explore the motion of fluid in gas-particle Sod test. Finally, we give a two-dimensional practical test case about plumbum plate ejecta motion under the high speed shock.
5.1 Particle transport tests
In this test, we investigate the motion of particles immersed in a fluid under different state which is an important problem in many applications.
5.1.1 Transport in a quiescent fluid
The motion of a particle acted upon by no force except drag is the most basic model. We compute this problem to investigate the particle trajectories under the drag influence. Suppose a nonzero initial speed for the particle at the middle of a nozzle filled with fluid and compare the velocity of particle and fluid. According to the Figure 3, the velocity of the particle is decreasing during the motion, with the deceleration gradually decreasing. At the same time, the speed of the fluid reaches the peak rapidly within a short period and then decreases. The speed difference between the particle and fluid ultimately converges, contrary to the intuition that the fluid velocity keeps increasing until they become consistent.
$t=0.2$
$t=0.8$
Figure 3: Drag model
Next based on the above tests, we further conduct the extend tests, investigating the speed variations by setting different radius for the particles in Figure 4. Obviously, as the particle radius increases, the speed decreases more slowly. A noteworthy phenomenon is that when the particle becomes sufficiently large, it is almost unaffected by the fluid, a characteristic reflected in subsequent numerical results.
Time-Velocity
Time-Position
Figure 4: Different radius comparison in a quiescent fluid
5.1.2 Motion under constant acceleration
In the second series of tests, according to the Figure 5 and 6, the different-size particles are given a nonzero initial velocity into a quiescent fluid with a constant acceleration. At the start of the motion, the fluid velocity is less than all the particles when the particles are enforced to decelerate, and then accelerate when the fluid velocity exceeds the particle velocity. Analyzing the curves of $r=0.0001$ and $r=0.001$, it can be observed that once the particle velocity is lower than that of the fluid, the particle accelerates with the same acceleration as the fluid, consistently lagging behind it. Similarly, when the particle becomes sufficiently large, for instance $r=1$, it is scarcely affected.
In addition, it is obvious that the particles eventually move with the different velocity differences from the fluid. Suppose that the acceleration of fluid is $\bm{a}_{g}$, the velocity difference is $\Delta\bm{v}=\bm{v}_{g}-\bm{v}_{p}$. There is the equation that $\bm{a}_{g}=\bm{F_{d}}/m_{p}=\bm{a}_{p}$ when the motion reaches the balance. In a general physical computation, drag force is positive correlation with the $\Delta\bm{v}$, we have $\bm{F_{d}}=\kappa\Delta\bm{v}$ in which $\kappa$ is a variable positive coefficient related to radius, density and viscosity when we used the different drag force model affecting the final velocity difference. For instance, in Stoke’s drag model, there are $\kappa=6\pi r_{p}\mu_{g}$ and $\Delta\bm{v}=\bm{a}_{g}r_{p}/6\mu_{g}$ which is consistent with the numerical results.
Time-Velocity
Figure 5: Different radius comparison in a constant acceleration fluid
Time-Position
Figure 6: Different radius comparison in a constant acceleration fluid
5.1.3 Motion under sinusoidal acceleration
In the final test, the different-size particles are given a nonzero initial velocity into a quiescent fluid with a sinusoidal acceleration, with the result in Figure 7. Comparing with the black line
Time-Velocity
Time-Position
Figure 7: Different radius comparison in a sinusoidal acceleration fluid
5.2 Multiphase Sod test
We are concerned with a shocktube problem[10] in the region $[-0.5,0.5]\times[0,0.1]$ of which solution consists of a left rarefaction, a contact and a right shock. We consider a shock tube of unity length. We test a series of initial conditions containing the particles of different size and analysis the influence of interaction to the flow. In the simulation, particles are uniformly distributed in the all cells of the region and the flow phase is an ideal gas. At the initial time, the states on the left-hand and right-hand sides of $x=0$ are constant. The left-hand state is a high pressure fluid characterized by $\left(\rho_{l},P_{l},u_{l}\right)=(9.6,1013250,0)$, and the right-hand state is a low pressure fluid defined by $\left(\rho_{r},P_{r},u_{r}\right)=(1.2,101325,0)$, which references the [32].
For the multiphase Sod shock tube, there is no exact solution. But we can analysis the physical interaction and results according to the solution of numerical simulation. In the numerical tests, we adjust the particles’ radius which achieve the different volume percentages, namely $0\%$, $0.01\%$, $4\%$, to research the effects on the fluid motion. Then in the table 1, we have the initial conditions for particles.
Table 1: Initial conditions for particles
Case volume percentage radius velocity
1 $0\%$ 00 0
2 $0.01\%$ $00010$
3 $4\%$ $0020$
In this tests, we run it in the 2-D region. But for the clearer comparison, we provide both 1-D and 2-D figure. Firstly in figure 8,9 and 10 there are the results for the motion of fluid without particles at time step $t=0.0004$.
Figure 8: 1-D and 2-D Density
Figure 9: 1-D and 2-D Velocity
Figure 10: 1-D and 2-D Pressure
Then we change the radius for the particles and give the results in figure 11,12 and 13. From the 1-D results, comparing the black line and red line represented the different the situation of particles. We can conclude that when the particle volume fraction is $0.01\%$, the influence on the fluid is almost negligible. It is difficult to discern especially in density and pressure curves, but a slight effect observed in the velocity curve. This suggests that under the above conditions, particles primarily affect the fluid velocity through force source terms.
Figure 11: 1-D and 2-D Pressure
Figure 12: 1-D and 2-D Pressure
Figure 13: 1-D and 2-D Pressure
Next as we increase the volume percentage of the particles to $4\%$, we can get a significant influence on fluid motion comparing pressure and density which is similar to the results in [32]. The results are as follows in 14:
Density
Pressure
Figure 14: Density and Pressure with $4\%$ particles
Remark 5.1.
Although the equations of our model are concerned with the heat exchange between the two phases, we usually ignore it by setting the same temperature in numerical tests. Actually apart from a few examples, energy exchange caused by temperature are not a concern.
5.3 Conclusion
Because of the interactions between two phases, the motion of fluid and particles is highly correlated. Under certain extreme and complex physical conditions, the velocity of particles or fluid could be too high to compute in the lagrangian form. We design the ALE method to solve the problem which does well in adapting to various situations and tracking particles and fluid.
The above numerical results can to some extent represent the mechanisms of particles influence on fluid. However, it doesn’t mean that the same results or influence can be obtained under any conditions. In numerical experiments, through a large amount of numerical tests, we found that the results are highly dependent on the choice fo fluid viscosity coefficient which indicates that the motion trend of fluid and particles is highly sensitive to the fluid viscosity coefficient. We need to analysis the realistic physical background when we do the numerical computations. Additionally, we didn’t give the change of internal energy in the above tests because the impact of thermal source term on the fluid is negligible compared with force source term in the case that there is almost no difference between the temperature of fluid and particles. Furthermore, due to the low volume fraction of particles, the fluid and particles have the same pressure at the same position of the region. As the fraction increases, it is necessary to consider the effect of pressure differences on motion in the computational process.
Acknowledgment
This research was supported by the National Natural Science Foundation of China (grant No. 12272059, 11871113, 22341302), the National Key Research and Development Program of China (grant No. 2020YFA0713602, 2023YFA1008803), and the Key Laboratory of Symbolic Computation and Knowledge Engineering of Ministry of Education of China housed at Jilin University.